# Direct Detection of Dark Photon Dark Matter with the James Webb Space Telescope

### v2

@@ Abstract
In this study, we propose an investigation into dark photon dark matter (DPDM) within the infrared frequency band, utilizing highly sensitive infrared light detectors commonly integrated into space telescopes, such as the James Webb Space Telescope (JWST). The presence of DPDM induces electron oscillations in both the reflectors and the interior of the detectors. Consequently, these oscillating electrons can emit monochromatic electromagnetic waves with a frequency almost equivalent to the mass of DPDM. By estimating the signal generated by DPDM inside the detector and comparing with observation data, we establish constraints on the kinetic mixing between the photon and dark photon within the range [10, 500] THz. Despite JWST not being optimized for DPDM searches, our findings reveal constraints comparable to those obtained from the XENON1T experiment in the laboratory, as well as astrophysical constraints from solar emission. Additionally, we propose to modify the configuration of JWST optical elements to focus the DPDM induced signal onto the detector. By employing the stationary phase approximation, we can demonstrate that when the size of the reflector significantly exceeds the wavelength of the electromagnetic wave, the contribution to the electromagnetic wave field at a given position primarily stems from the surface unit perpendicular to the relative position vector. This simplification results in the reduction of electromagnetic wave calculations to ray optics. We show that by rearranging the position of reflectors, JWST can achieve a sensitivity stronger than the existing limits by 1 or 2 orders of magnitude.
@@ 1 Introduction
Due to the vast range of unknown mass of the dark photon dark matter (DPDM), there are various detecting methods accordingly [31, 32]. The relevant searches for dark photon DM are haloscope experiments [33, 34, 35, 36, 37, 38], dish antenna experiments [39, 40, 41, 42], plasma telescopes [43], CMB spectrum distortion [9, 44] and radio telescopes [45, 46, 47]. The searches include direct detection of local DPDM in laboratories and observation on its impact in the early universe.
Recently, we proposed to search for Dark Photon Dark Matter (DPDM) conversion, specifically $A^{\prime}\to\gamma$, locally using radio telescopes such as FAST and LOFAR [46]. For instance, the FAST radio telescope, equipped with a large dish antenna, converts DPDM into a regular photon field. In each small surface area of FAST, an oscillating electric dipole is generated by the DPDM field, with a frequency matching the DPDM mass. Summing up the contributions from each surface area yields the total converted electromagnetic field. The original proposal utilized a spherical reflector, causing the electromagnetic field to constructively focus on the spherical center [39, 48, 41]. This concept has been previously employed in shielded room-sized experiments by various works, utilizing variations such as plane/parabolic reflectors or dipole antennas [49, 50, 42, 51, 52, 53, 54, 55].
In this article, we would like to explore this idea with the recent new telescope, the James Webber Science Telescope (JWST), which is running in space. JWST covers the frequency range of 10–500 THz for infrared astronomy. Our searches for DPDM benefit from JWST’s high frequency resolution $R=f/\Delta f$, which ranges from 4 to 3000, depending on different obervation modes [56]. Also, since it works in space, we expect it to have a much lower noise background than terrestrial facilities. The Near Infrared Spectrograph (NISpec) and the Mid-Infrared Instrument (MIRI), two instruments carried by JWST, are especially useful for searching for DPDM. By analyzing data collected by the two instruments, assuming DPDM directly converts to monochromatic electromagnetic signal at their pixel detectors, we get the upper limits for the DPDM-photon coupling constant, $\epsilon\sim 10^{-10}-10^{-12}$ in the frequency range $10-500$ THz at $95\%$ confidence level (C.L.). Additionally, we propose that by rearranging the position of the optical elements of JWST, a stronger constraint $\epsilon\sim 10^{-12}-10^{-14}$ can be achieved.
@@
This work is structured as follows: In Section I, we present the background of DPDM conversion. In Section II, we offer a mathematical proof of the reduction from complex electroweak wave calculation to ray optics using the stationary phase approximation. Section III gives an introduction to the optical telescope element (OTE) of JWST. In section IV, we set a constraint on the kinetic mixing parameter $\epsilon$ by analyzing JWST observation data. Section V employs the idea of modifying the configuration of JWST optical elements, projecting a limit on the kinetic mixing parameter for future experiments. Finally, in Section VI, we draw our conclusions.
@@ 2 High Frequency Approximation
In order to calculate the EM signals induced by DPDM on a metal reflector plate, the most direct way is to divide the reflector into many small patches and sum up the induced EM signals over all of them. The length of each small patch is required to be much smaller than the wavelength $\lambda$ of the induced EW signal while at the same time much larger than the reflector thickness. Consequently, to ensure enough accuracy, the simulation mesh must be fine enough for the distance between mesh points to be smaller than the wavelength. This method works well in the case of FAST telescope [57]. The FAST detects radio photons around $1$ GHz and has a reflector roughly $500$ m in diameter. Therefore, we only need to divide the FAST reflector into $\sim 10^{6}$ patches for an accurate simulation. On the other hand, JWST works at a much higher frequency range, $10-500$ THz (i.e., the photon wavelength $\lambda\sim 0.6-30\leavevmode\nobreak\ \mathrm{\mu m}$), and the diameter of the JWST’s primary mirror is $D=6.6$ meter. To achieve an acceptable level of accuracy in simulating the JWST case, we require over $10^{15}$ patches, which is finer than the FAST case by orders of magnitude. However, such a fine mesh imposes an immense computational burden, rendering it unfeasible to simulate signal strength using available computer resources.
@@
Fortunately, as we will see, we can theoretically demonstrate that calculating the induced EM signals becomes considerably more straightforward in the high-frequency regime, $D\gg\lambda$. Similar to the case of reaching the ray optics at the high-frequency limit of the wave optics, we can establish that the strength of the DPDM-induced signal can be obtained using a ray-optical method. This simplification ultimately leads to a set of algebraic equations that can be easily calculated, even manually without the need of a fine-mesh simulation. The physical interpretation of such a simplification is that, in the high-frequency regime, interferences between different patches on the reflector have a negligible impact on the final result.
@@
To be more specific, this simplification works primarily due to two key factors. Firstly, the phases of the electric fields contributed by different patches on a plate vary significantly, while their strengths remain relatively the same. This results in significant cancellations between the electric fields generated by different patches. Secondly, one significant parameter is the coherence length of DPDM. The DPDM and the induced photon have the same energy, but the coherent length of DPDM, $\lambda^{\prime}$, is much larger than the wavelengh of the induced photons, $\lambda$. This is due to the non-relativistic nature of dark matter with a low speed $v_{\rm DM}\sim 10^{-3}$, which gives $\lambda^{\prime}\sim\lambda/v_{\rm DM}$. Importantly, $\lambda^{\prime}$ is still significantly smaller than the diameter of the JWST’s mirrors. As a result, the interferences between different patches are dampened by incoherence.
@@ 2.1 Monochromatic DPDM
$_{t}$ here means the average over time. In principle, for any reflectors, we can numerically simulate the DPDM-induced EM waves using these formulas. However, as we discussed above, such a simulation requires a very fine mesh that is hard to realize in computers for the JWST case. As we are going to see below, we figure out a more analytical method applicable in the high-frequency regime.
@@
$\frac{\partial}{\partial u}|\bm{r}-\bm{r^{\prime}}(u,v)|=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\partial}{\partial v}|\bm{r}-\bm{r^{\prime}}(u,v)|=0,$ (2.10)
@@
$(\bm{r}-\bm{r^{\prime}})\cdot\frac{\partial\bm{r}^{\prime}}{\partial u}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (\bm{r}-\bm{r^{\prime}})\cdot\frac{\partial\bm{r}^{\prime}}{\partial v}=0.$ (2.11)
$u$ and $v$ can be understood as two parameters describing a certain patch of the surface, so $\partial\bm{r}^{\prime}/\partial u$ and $\partial\bm{r}^{\prime}/\partial v$ are tangent vectors at $\bm{r}^{\prime}$. Therefore, Eq. (2.11) tells us that $\bm{r}-\bm{r}^{\prime}$ is perpendicular to the tangent plane at $\bm{r}^{\prime}$. This result can be interpreted in a more intuitive way. Considering a conductor surface, at each point $\bm{r}^{\prime}$ on the surface, a light ray is only emitted in the normal direction, then the signal received at the position $\bm{r}$ is the sum of the light rays passing through the position $\bm{r}$. Therefore, we see the calculations in the JWST case can be accomplished within the framework of ray optics. The key difference with conventional ray optics is that a light ray induced by DPDM is always perpendicular to the local surface from which it is emitted, regardless of the direction in which the DPDM is incident.
@@
where $\hat{\bm{n}}(\hat{\bm{r}}_{j})$ is the out-pointing normal direction of the conductor surface at $\hat{\bm{r}}_{j}$. The interference term comes from the cross products of contributions from $\hat{\bm{r}}_{j}$ and $\hat{\bm{r}}_{k}$, with $j\neq k$. If there is only one $\hat{\bm{r}}_{j}$, there is only one term in the summation and the interference term drops out. As a consistency check, one can compare (2.13) with the result of infinitely large metal plate given in appendix I-A of Ref. [57].
@@ 2.2 Non-monochromatic DPDM
According to the Standard Halo Model, dark matter has a truncated Maxwellian distribution in the momentum space. Consequently, we should take into consideration the effect of finite coherence length of DPDM. The dark photon field at a location $\bm{x}$ can be expressed as,
@@
[…]
@@
×$\times[\bm{\tau}(\bm{r^{\prime\prime}})\times(\bm{r}-\bm{r^{\prime\prime}})]\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)+i\bm{k}^{\prime}\cdot(\bm{r^{\prime}}-\bm{r^{\prime\prime}})}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right).$ (2.18)
This expression can be simplified after some computation, the detail is given in appendix A.
@@
×$\times e^{-\frac{1}{8}k_{0}^{2}|\bm{r^{\prime}}-\bm{r^{\prime\prime}}|^{2}}\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right)$ (2.19)
If we choose the reflector to be spherical, some analytic results can be derived. In the $k_{0}\rightarrow 0$ limit, i.e. the infinite correlation length limit, the result is
@@
where $\gamma$ is the angle between the polarization vector $\bm{n}_{0}$ and z-direction, $\theta_{0}$ describes how large the spherical surface is, with $\theta_{0}=0$ for the surface shrinking to a point and $\theta_{0}=\pi$ for the surface becoming a full sphere. This is the same as what we obtained in [57]. We are also interested in the $k_{0}\rightarrow\infty$ limit, i.e. the zero correlation length limit, the result is
@@
Interestingly, the flux in the high frequency limit doesn’t depend on the radius of the sphere, but it does depend on $\theta_{0}$. Note that (2.21) only applies when the radius of the reflector is much larger than the dark photon wavelength, which is $10^{3}$ times the same-frequency EM wavelength. Naively, larger reflectors produce stronger signal, but (2.21) tells us that the signal saturates when the reflector is much larger than the dark photon wavelength.
@@
$\bm{r}-\bm{r}^{\prime}\perp\mbox{tangent plane at }\bm{r}^{\prime},\quad\bm{r}^{\prime}=\bm{r}^{\prime\prime}$ (2.22)
@@
Again, $\hat{\bm{n}}(\hat{\bm{r}}_{i})$ is the out-directed normal vector at $\hat{\bm{r}}_{i}$, and $\hat{\bm{A^{\prime}}}(\hat{\bm{r}}_{i})=|\bm{A}^{\prime}|\bm{\tau}(\hat{\bm{r}}_{i})$. Note that if the set of solutions to (2.22) is not discrete, we should change the sum into an integration. If one wishes to average over all possible polarization, the result is
@@
In the JWST case, all approximation conditions applied in this section are satisfied, so one may directly use (2.24) to compute the dark photon flux density.
3 The Optical Telescope Element (OTE) of JWST
The Optical Telescope Element (OTE) of the James Webb Space Telescope (JWST) comprises a primary mirror, a secondary mirror, a tertiary mirror, and a fine steering mirror. A sketch of the mirror system is shown in Figure 1. Detailed parameters for these optical components are available in JWST documentation [61], and we have also summarized them in Table 1.

### v3

@@ Abstract
In this study, we propose an investigation into dark photon dark matter (DPDM) within the infrared frequency band, utilizing highly sensitive infrared light detectors commonly integrated into space telescopes, such as the James Webb Space Telescope (JWST). The presence of DPDM induces electron oscillations in both the reflectors and the interior of the detectors. Consequently, these oscillating electrons can emit monochromatic electromagnetic waves with a frequency almost equivalent to the mass of DPDM. By employing the stationary phase approximation, we can demonstrate that when the size of the reflector significantly exceeds the wavelength of the electromagnetic wave, the contribution to the electromagnetic wave field at a given position primarily stems from the surface unit perpendicular to the relative position vector. This simplification results in the reduction of electromagnetic wave calculations to ray optics. Through a careful analysis of photon generation induced by DPDM on the various optical elements of JWST, we find that the contribution of these photons to the detected signal is negligible. Nevertheless, we propose a modified configuration of the JWST mirrors that would enable the DPDM-induced photons to be focused onto the detector. This approach can be applied to future space telescopes during their ground-testing phases. Using the JWST parameters as a representative example, the achievable upper limits on the DPDM-photon mixing constant are $\epsilon\sim 10^{-12}-10^{-14}$ in the frequency range 1050010-500 THz at the 95% confidence level. This reveals the strong potential of future space telescopes for DPDM detection during ground testing, with sensitivities exceeding current limits by 1 to 2 orders of magnitude compared with the XENON1T result and the solar cooling bound.
@@ 1 Introduction
Due to the vast range of unknown mass of the dark photon dark matter (DPDM), there are various detecting methods accordingly [33, 22]. The relevant searches for DPDM are haloscope experiments [27, 76, 42, 16, 15, 43], dish antenna experiments [45, 47, 46, 52], plasma telescopes [38], CMB spectrum distortion [14, 59], and radio telescopes [11, 8, 10]. The searches include direct detection of local DPDM in laboratories and observation of its impact in the early universe.
Recently, we proposed to search for the DPDM conversion, specifically $A^{\prime}\to\gamma$, locally using radio telescopes such as FAST and LOFAR [8]. For instance, the FAST radio telescope, equipped with a large dish antenna, converts DPDM into a regular photon field. In each small surface area of FAST, an oscillating electric dipole is generated by the DPDM field, with a frequency matching the DPDM mass. Summing up the contributions from each surface area yields the total converted electromagnetic field. The original proposal utilized a spherical reflector, causing the electromagnetic field to constructively focus on the spherical center [45, 48, 46]. This concept has been previously employed in shielded room-sized experiments by various works, utilizing variations such as plane/parabolic reflectors or dipole antennas [73, 72, 52, 75, 39, 21, 12, 17].
In this article, we would like to explore this idea with the recent new telescope, the James Webb Space Telescope (JWST), which is running in space. JWST covers the frequency range of 10–500 THz for infrared astronomy. Our searches for DPDM benefit from JWST’s high frequency resolution $R=f/\Delta f$, which ranges from 4 to 3000, depending on different observation modes [36]. The Near Infrared Spectrograph (NISpec) and the Mid-Infrared Instrument (MIRI), two instruments carried by JWST, are especially useful for searching for DPDM. We carefully analyze the photon signals induced by DPDM in the various optical components of JWST and find that their contribution to the final detected signal is negligible. This is primarily because JWST is optimized for observing distant astrophysical sources and is not ideally suited for detecting DPDM-induced photons, which propagate along directions normal to the local mirror surfaces. However, with appropriate adjustments to the mirror positioning during the pre-launch ground-testing phase, we show that space telescopes such as JWST can become powerful instruments for probing DPDM. This approach can be extended to future missions such as LUVOIR [74], HabEx [37], Origins [60], Roman (WFIRST) [71], etc. Using the parameters of NIRSpec and MIRI onboard JWST as representative examples, we obtain upper limits on the DPDM–photon kinetic mixing parameter of $\epsilon\sim 10^{-12}-10^{-14}$ in the frequency range 1050010-500 THz at $95\%$ confidence level (C.L.). These results highlight the strong potential of future space telescopes for DPDM searches, provided their mirror configurations can be moderately adjusted during ground-based testing.
@@
This work is structured as follows: In section 2, we provide a mathematical proof demonstrating how the complex electromagnetic-wave formulation reduces to ray optics under the stationary phase approximation. In section 3, we carefully analyze the optical system of JWST, including the Optical Telescope Element and the detector optics, and show that the contribution of photons induced by DPDM on the various optical components of JWST to the detected signal is negligible. Then, in section 4, we demonstrate that, with appropriate adjustments to the positioning of the telescope’s mirrors during the ground-testing phase, the telescope would be able to achieve significant sensitivity to DPDM detection – an approach that could be applied to upcoming space telescope missions. Finally, section 5 summarizes our results.
@@ 2 High Frequency Approximation
In order to calculate the EM signals induced by DPDM on a metal reflector plate, the most direct way is to divide the reflector into many small patches and sum up the induced EM signals over all of them. The length of each small patch is required to be much smaller than the wavelength $\lambda$ of the induced EW signal while at the same time much larger than the reflector thickness. Consequently, to ensure enough accuracy, the simulation mesh must be fine enough for the distance between mesh points to be smaller than the wavelength. This method works well in the case of FAST telescope [9]. The FAST detects radio photons around $1$ GHz and has a reflector roughly $500$ m in diameter. Therefore, we only need to divide the FAST reflector into $\sim 10^{6}$ patches for an accurate simulation. On the other hand, JWST works at a much higher frequency range, 1050010-500 THz (i.e., the photon wavelength $\lambda\sim 0.6-30\penalty 10000\ \mathrm{\mu m}$), and the diameter of the JWST’s primary mirror is $D=6.6$ meters. To achieve an acceptable level of accuracy in simulating the JWST case, we require over $10^{15}$ patches, which is finer than the FAST case by orders of magnitude. However, such a fine mesh imposes an immense computational burden, rendering it unfeasible to simulate signal strength using available computer resources.
@@
Fortunately, as we will see, we can theoretically demonstrate that calculating the induced EM signals becomes considerably more straightforward in the high-frequency regime, $D\gg\lambda$. Similar to the case of reaching the ray optics at the high-frequency limit of the wave optics, we can establish that the strength of the DPDM-induced signal can be obtained using a ray-optical method. This simplification ultimately leads to a set of algebraic equations that can be easily calculated, even manually without the need for a fine-mesh simulation. The physical interpretation of such a simplification is that, in the high-frequency regime, interferences between different patches on the reflector have a negligible impact on the final result.
@@
To be more specific, this simplification works primarily due to two key factors. Firstly, the phases of the electric fields contributed by different patches on a plate vary significantly, while their strengths remain relatively the same. This results in significant cancellations between the electric fields generated by different patches. Secondly, one significant parameter is the coherence length of DPDM. The DPDM and the induced photon have the same energy, but the coherent length of DPDM, $\lambda^{\prime}$, is much larger than the wavelength of the induced photons, $\lambda$. This is due to the non-relativistic nature of dark matter with a low speed $v_{\rm DM}\sim 10^{-3}$, which gives $\lambda^{\prime}\sim\lambda/v_{\rm DM}$. Importantly, $\lambda^{\prime}$ is still significantly smaller than the diameter of the JWST’s mirrors. As a result, the interferences between different patches are dampened by incoherence.
@@ 2.1 Monochromatic DPDM
$\left<...\right>_{t}$ here means the average over time. In principle, for any reflectors, we can numerically simulate the DPDM-induced EM waves using these formulas. However, as we discussed above, such a simulation requires a very fine mesh that is hard to realize in computers for the JWST case. As we are going to see below, we figure out a more analytical method applicable in the high-frequency regime.
@@
$\frac{\partial}{\partial u}|\bm{r}-\bm{r^{\prime}}(u,v)|=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \frac{\partial}{\partial v}|\bm{r}-\bm{r^{\prime}}(u,v)|=0,$ (2.10)
@@
$(\bm{r}-\bm{r^{\prime}})\cdot\frac{\partial\bm{r}^{\prime}}{\partial u}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ (\bm{r}-\bm{r^{\prime}})\cdot\frac{\partial\bm{r}^{\prime}}{\partial v}=0.$ (2.11)
$u$ and $v$ can be understood as two parameters describing a certain patch of the surface, so $\partial\bm{r}^{\prime}/\partial u$ and $\partial\bm{r}^{\prime}/\partial v$ are tangent vectors at $\bm{r}^{\prime}$. Therefore, Eq. (2.11) tells us that $\bm{r}-\bm{r}^{\prime}$ is perpendicular to the tangent plane at $\bm{r}^{\prime}$. This result can be interpreted in a more intuitive way. Considering a conductor surface, at each point $\bm{r}^{\prime}$ on the surface, a light ray is only emitted in the normal direction, then the signal received at the position $\bm{r}$ is the sum of the light rays passing through the position $\bm{r}$. Therefore, we see that the calculations in the JWST case can be accomplished within the framework of ray optics. The key difference with conventional ray optics is that a light ray induced by DPDM is always perpendicular to the local surface from which it is emitted, regardless of the direction in which the DPDM is incident.
@@
where $\hat{\bm{n}}(\hat{\bm{r}}_{j})$ is the out-pointing normal direction of the conductor surface at $\hat{\bm{r}}_{j}$. The interference term comes from the cross products of contributions from $\hat{\bm{r}}_{j}$ and $\hat{\bm{r}}_{k}$, with $j\neq k$. If there is only one $\hat{\bm{r}}_{j}$, there is only one term in the summation, and the interference term drops out. As a consistency check, one can compare (2.13) with the result of an infinitely large metal plate given in Appendix I-A of Ref. [9].
@@ 2.2 Non-monochromatic DPDM
According to the Standard Halo Model, dark matter has a truncated Maxwellian distribution in momentum space. Consequently, we should take into consideration the effect of finite coherence length of DPDM. The dark photon field at a location $\bm{x}$ can be expressed as,
@@
where $\theta(\bm{k}^{\prime})$ is a random phase associated with the $\bm{k}^{\prime}$ mode and $\omega=\sqrt{\bm{k}^{\prime 2}+m_{A^{\prime}}^{2}}$ is the energy. $b$ is a normalization factor. We have $k_{0}=m_{A^{\prime}}v_{0}$ and $k_{\mathrm{esc}}=m_{A^{\prime}}v_{\mathrm{esc}}$, where $v_{0}\approx$235\text{\,}\mathrm{k}\mathrm{m}\mathrm{/}\mathrm{s}$\$ is the most probable velocity and $v_{\mathrm{esc}}$ is the escape velocity of leaving the Galaxy at the position of the solar system which is about $500\penalty 10000\ {\rm km/s}$ [30, 32]. Due to randomness, we assume that there is no correlation between different momentum modes,
@@
$\times[\bm{\tau}(\bm{r^{\prime\prime}})\times(\bm{r}-\bm{r^{\prime\prime}})]\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)+i\bm{k}^{\prime}\cdot(\bm{r^{\prime}}-\bm{r^{\prime\prime}})}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right).$ (2.18)
This expression can be simplified after some computation; the detail is given in Appendix A.
@@
$\times e^{-\frac{1}{8}k_{0}^{2}|\bm{r^{\prime}}-\bm{r^{\prime\prime}}|^{2}}\mathrm{Re}\left(\frac{e^{im_{A^{\prime}}(|\bm{r}-\bm{r^{\prime}}|-|\bm{r}-\bm{r^{\prime\prime}}|)}}{|\bm{r}-\bm{r^{\prime}}|^{3}|\bm{r}-\bm{r^{\prime\prime}}|^{2}}\right)$ (2.19)
If we choose the reflector to be spherical, some analytic results can be derived. In the $k_{0}\rightarrow 0$ limit, i.e., the infinite correlation length limit, the result is
@@
where $\gamma$ is the angle between the polarization vector $\bm{n}_{0}$ and z-direction, $\theta_{0}$ describes how large the spherical surface is, with $\theta_{0}=0$ for the surface shrinking to a point and $\theta_{0}=\pi$ for the surface becoming a full sphere. This is the same as what we obtained in [9]. We are also interested in the $k_{0}\rightarrow\infty$ limit, i.e., the zero correlation length limit, the result is
@@
Interestingly, the flux in the high frequency limit doesn’t depend on the radius of the sphere, but it does depend on $\theta_{0}$. Note that (2.21) only applies when the radius of the reflector is much larger than the dark photon wavelength, which is $10^{3}$ times the same-frequency EM wavelength. Naively, larger reflectors produce a stronger signal, but (2.21) tells us that the signal saturates when the reflector is much larger than the dark photon wavelength.
@@
[…]
@@
Again, $\hat{\bm{n}}(\hat{\bm{r}}_{i})$ is the out-directed normal vector at $\hat{\bm{r}}_{i}$, and $\hat{\bm{A^{\prime}}}(\hat{\bm{r}}_{i})=|\bm{A}^{\prime}|\bm{\tau}(\hat{\bm{r}}_{i})$. Note that if the set of solutions to (2.22) is not discrete, we should change the sum into an integration. If one wishes to average over all possible polarizations, the result is
@@
For a space telescope such as JWST, all approximation conditions applied in this section are satisfied, so one may directly use (2.24) to compute the dark photon flux density. This will finally be applied in section 4 for a modified JWST configuration.
3 The Optical System of JWST
The optical system of JWST can be separated into two parts, the Optical Telescope Element (OTE) and the detector optics. OTE is followed by the detector optics. DPDM may generate signals in both OTE and the detector optics. Below, we discuss these two cases and demonstrate why the contribution of the DPDM-induced photons to the final detected signal is weak.